[Koji, Steve, Jamie, Jenne, Yuta]
We opened BS and ITMX chambers, took lots of photos, and closed them with heavy doors.
I turned off high voltage power supplies for PZTs and blocked PSL beam. We are ready for the pumping tomorrow.
Important photos we took:
- positions of green optics at BS chamber, which was moved on the vent on Aug 2011
- positions of PZT mirrors and cable connectors at BS chamber, which will be replaced with tip-tilts on the next vent
- arrow on PR2 pointing HR (it was correct)
- tried to take photos of clipping IR beam at BS OSEM holder from ITMX chamber
We also took bunch of other photos.
Beam dump needed at BS chamber:
We also checked some un-dumped beams at BS chamber. We need dumps;
- behind MMT1, for unwanted transmitted beam
- behind IPPOSSM3, for unwanted transmitted beam (IPPOSSM3 is the last mirror in BS chamber for IPPOS)